OpenStudy (superfly123):

If the legs of a right triangle are x-3 and 9, and its hypotenuse is x, find the length of its smallest side.

OpenStudy (superfly123):

I set up ( x-3 ) ^2 + 9^2 = x^2

jhonyy9 (jhonyy9):

so this is a right triangle - do you know Pythagora's theorem ?

OpenStudy (superfly123):

Yes

jhonyy9 (jhonyy9):

you need just use it and will get what you need

OpenStudy (superfly123):

Yeh, but i dont know how set up everything

jhonyy9 (jhonyy9):

so what you have wrote above - this equation just you need to solve it for x

OpenStudy (superfly123):

okay so x^2 -6x +9 + 81 = x^2 x^2 -6x + 90 = x^2 -6x = 90 so i get -15

OpenStudy (superfly123):

But all my friends say positive 15 i dont know how though

jhonyy9 (jhonyy9):

there above you have wrote (x-3)^2 +9^2 = x^2 yes ?

OpenStudy (superfly123):

yes

jhonyy9 (jhonyy9):

but check it please that there is something wrong - after you eliminated the x^2 what will remain there ?

OpenStudy (superfly123):

-6x and 90

jhonyy9 (jhonyy9):

yes but hence -6x+90 = 0 so what will get the subtract from both sides 90 ?

OpenStudy (superfly123):

oh -90

jhonyy9 (jhonyy9):

yes sure

OpenStudy (superfly123):

resulting in 15 from there

OpenStudy (superfly123):

so the smallest side is 9

jhonyy9 (jhonyy9):

yes exactly - do you see it now your friends was rights

OpenStudy (superfly123):

oh ok yes, so the smallest side was already given

jhonyy9 (jhonyy9):

yes - this is ok. now ? do you understand it all sure ?

OpenStudy (superfly123):

Yup, thanks

jhonyy9 (jhonyy9):

np welcome was my pleasure anytime ---- good luck bye